OpenStudy (anonymous):

The perimeter of a rectangle is 56 inches. The length is 4 more than 3 times the width. What are the dimensions of the rectangle? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

Write the system of equations that would be used to solve this problem. 2L +2W=56 L= 3W+4 Solve the system of equations in problem #3. ANSWER: Solve by substitution: 2(3W+4) +2W = 56 6W+8 +2W = 56 8W = 56-8 8W=48 W=6 From L=3W+4 L=3(6)+4 =18+4 =22 What is the length of one side of the rectangle? 6 inches 12 inches 22 inches 44 inches ANSWER: This question is ambiguous because there are four sides. If the questioner means the width, then it’s 6 inches if the length, then it’s 22 inches.

OpenStudy (anonymous):

It asks for dimensions it's not ambiguous Your answer is correct the rectangle is 6 by 22

OpenStudy (anonymous):

yes :D